OpenStudy (anonymous):

You pay 3.3 cents per minute for long-distance phone calls on your phone. Your phone bill is $3.40. Write and solve an equation to find the number of minutes m of long-distance calls you made

OpenStudy (anonymous):

3.40 ÷ 3.3 = x

OpenStudy (anonymous):

0.033m=3.40 m=3.40/0.033 m=103.03 minutes